# Ratio Covers of Convex Sets and Optimal Mixture Density Estimation

### v1

@@ 1.5 Structure of the paper
In Section˜2 we study model aggregation over a finite family of densities and prove Theorem˜2. In Section˜3, we establish a sharp local covering bound in Hellinger distance for mixtures (Theorem˜3). Section˜4 is devoted to ratio covers, including the proof of Theorem˜1 and additional results on approximate Pareto curves. These tools are combined in Section˜5 to obtain the high-probability bound for convex aggregation (Theorem˜6). Further applications of ratio covers are presented in Section˜6: expected KL guarantees via a refined Yang–Barron argument (Section˜6.1) and KL covering numbers (Section˜6.2). We conclude with the misspecified setting in Section˜7.
@@ 2 Model aggregation
In this section, we provide a complete analysis of model aggregation over $M$ individual densities, where no assumptions are made on the individual densities. Despite its general novelty, this result is technically less involved and can be seen as a preliminary building block for a more complicated case involving the convex hull of arbitrary densities. In particular, even at the geometry level, we might not need to use the Birgé-Le Cam tournaments, but rather exploit simple properties of the MLE in this case.
@@
This bound is optimal in the sense that there are $M$ distributions such that no estimator can perform better up to multiplicative constant factors as shown in Mourtada (2025); van der Hoeven et al. (2025).
@@
  • Use the first $n$ observations to construct a data-dependent set of almost empirical risk minimizers
    $\mathcal{W}=\left\{p\in\mathcal{P}:\widehat{L}_{n}(p)-\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant\frac{2\log(4M/\delta)}{n}\right\}.$
@@
    $\mathcal{W}^{\prime}=\left\{\frac{1}{2}p+\frac{1}{2|\mathcal{W}|}\sum\limits_{q\in\mathcal{W}}q:p\in\mathcal{W}\right\}.$ (4)
  • Using the second half of the sample, run a version of the progressive density estimator of Lemma˜2 over $\mathcal{W}^{\prime}$ to output the final density $\widehat{p}$. Finally, if $n\leqslant 75\log(2/\delta)$ output instead $\widehat{p}=\frac{1}{M}\sum\limits_{p\in\mathcal{P}}p$.
Let us quickly explain why the procedure works. We first show that the construction of the set $\mathcal{W}$ of almost empirical likelihood maximizers provides a set of small Hellinger radius that contains $p^{\star}$. However, closeness in Hellinger radius does not imply closeness in the Kullback–Leibler divergence. This is where the smoothed class $\mathcal{W}^{\prime}$ helps: the construction does not inflate the Hellinger distance, but it allows one to control the density ratio. However, as an inevitable price of smoothing $p^{\star}$ does not have to belong to $\mathcal{W}^{\prime}$. This is where now the regret-based online-to-batch conversion techniques help, since they provide both the optimal tail bounds and allow to bypass the problem $p^{\star}\notin\mathcal{W}^{\prime}$. Remarkably, the final estimator is optimal up to multiplicative constant factors.
@@
$\mathcal{W}=\bigg\{p\in\mathcal{P}:\ \widehat{L}_{n}(p)-\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant\frac{2\log(2M/\delta)}{n}\bigg\}.$
Then, with probability at least $1-\delta$, one has $p^{\star}\in\mathcal{W}$ and
$\max_{p\in\mathcal{W}}\operatorname{H}^{2}(p^{\star},p)\leqslant\frac{2\log(2M/\delta)}{n}.$
@@
$\mathbf{P}\left(\widehat{L}_{n}(p)-\widehat{L}_{n}(p^{\star})\leqslant t\right)$ $=\mathbf{P}\left(\sqrt{\frac{\prod_{i=1}^{n}p(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}}\geqslant\exp(-nt/2)\right)$
@@
$\mathbf{P}\left(\exists p\in\mathcal{P}_{t}:\ \widehat{L}_{n}(p)-\widehat{L}_{n}(p^{\star})\leqslant t\right)\leqslant\sum_{p\in\mathcal{P}_{t}}\exp(-nt/2)\leqslant M\,\exp(-nt/2).$
Note that for any $p\in\mathcal{P}$, $\widehat{L}_{n}(p)-\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant t$ implies $\widehat{L}_{n}(p)-\widehat{L}_{n}(p^{\star})\leqslant t$, since $\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant\widehat{L}_{n}(p^{\star})$. Therefore,
$\mathbf{P}\left(\exists p\in\mathcal{P}:\widehat{L}_{n}(p)-\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant t\ \text{and}\ \operatorname{H}^{2}(p^{\star},p)\geqslant t\right)\leqslant M\exp(-nt/2).$
@@
$\mathbf{P}\left(\widehat{L}_{n}(p)-\widehat{L}_{n}(p^{\star})\leqslant-t\right)$ $=\mathbf{P}\left(\sqrt{\frac{\prod_{i=1}^{n}p(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}}\geqslant\exp(nt/2)\right)$
@@
$\mathbf{P}\left(\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant\widehat{L}_{n}(p^{\star})-t\right)\leqslant\sum_{p\in\mathcal{P}}\mathbf{P}\left(\widehat{L}_{n}(p)-\widehat{L}_{n}(p^{\star})\leqslant-t\right)\leqslant M\exp(-nt/2).$
@@
$\mathbf{P}\left(\widehat{L}_{n}(p^{\star})-\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)>t\right)\leqslant M\exp(-nt/2).$
Choose $t=\frac{2\log(2M/\delta)}{n}$. Then each of the two bad events above has probability at most $\delta/2$, so by a union bound they fail simultaneously with probability at most $\delta$. Hence, with probability at least $1-\delta$, every $p\in\mathcal{W}$ satisfies $\operatorname{H}^{2}(p^{\star},p)<t$ and also $\widehat{L}_{n}(p^{\star})-\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant t$, that is, $p^{\star}\in\mathcal{W}$. This yields the claim. ∎
@@
$\mathcal{W}=\left\{p\in\mathcal{P}:\widehat{L}_{n}(p)-\min_{q\in\mathcal{P}}\widehat{L}_{n}(q)\leqslant\frac{2\log(4M/\delta)}{n}\right\}$
@@
$\sup_{p,q\in\mathcal{W}^{\prime}}\sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\leqslant\log\bigl(1+|\mathcal{W}|\bigr)\leqslant\log(2M).$
@@
$\operatorname{KL}({p^{\star}},{\widehat{p}})\leqslant\operatorname{KL}({p^{\star}},{\widetilde{p}})+\frac{2\log(M)+\frac{25}{4}\left(e-2\right)\left(\log(2M)+1\right)\log\left(\frac{2}{\delta}\right)}{n}.$ (5)
Finally, we have by Lemma 1 and the convexity of the Hellinger distance squared, observing that the ratio of $p^{\star}$ and $\frac{1}{2}p^{\star}+\frac{1}{2|\mathcal{W}|}\sum\limits_{q\in\mathcal{W}}q$ is bounded by $2$,
@@
$\leqslant\frac{1}{(\sqrt{e}-1)^{2}}\operatorname{H}^{2}\left(p^{\star},\frac{1}{|\mathcal{W}|}\sum\limits_{q\in\mathcal{W}}q\right)\leqslant\frac{2\log(4M/\delta)}{(\sqrt{e}-1)^{2}n}.$
@@
Note that we used a sample of size $2n$; for any original sample size at least $2$, splitting into two equal parts (dropping at most one observation) increases the absolute constant by at most a factor of $3$ (the worst case is a total sample size of $3$). Finally, if $n\leqslant 75\log(2/\delta)$ we output $\widehat{p}=\frac{1}{M}\sum\limits_{p\in\mathcal{P}}p$, which trivially gives the upper bound $\log(M)$. The claim follows.∎
@@ 4 Ratio covers of convex sets
After approximating the target distribution in Hellinger distance (as we did above, for instance in the special case of finite mixtures), a central question in this work is how to convert this control into a small set of candidate distributions such that at least one is close to $p^{\star}$ in KL divergence. We recall from Lemma 1 that the KL divergence $\operatorname{KL}({p},{q})$ can be bounded in terms of the Hellinger distance together with a factor depending on the maximal likelihood ratio $\sup_{x\in\mathcal{X}}\frac{p(x)}{q(x)}$. In particular, if $p$ and $q$ lie in the same squared Hellinger ball of radius $\varepsilon$ and the ratio $\sup_{x\in\mathcal{X}}\frac{p(x)}{q(x)}$ is bounded by a constant, then one obtains a bound of the form $\operatorname{KL}({p},{q})\leqslant C\,\varepsilon$ for some constant $C>0$ depending only on that ratio bound. This motivates the notion of a ratio cover, which we will later apply to Hellinger balls to obtain the KL covers needed in our applications.
@@
Our goal is to show the existence of $C_{1}$-ratio covers of size $(C_{2})^{d}$ for convex sets $K$, with constants $C_{1},C_{2}\geqslant 1$. In our proof of Theorem 6 for mixture models, we will ultimately pay a term proportional to $\log(|A|)/n$ for a ratio cover of size $|A|$: hence even a ratio cover of size $|A|=(\log d)^{d}$ would yield a suboptimal upper bound, and similarly any super-constant ratio $\alpha$ would yield suboptimal results (both issues are also suboptimal for our other applications). With this motivation, we state one of our main results.
See 1
@@
First, the constraint $A\subset K$ is essential: without it, there is a trivial “cover” obtained by choosing the single vector $\phi\in\mathbb{R}_{+}^{d}$ with coordinates $\phi_{j}=\max_{\theta\in K}\theta_{j}$, which need not belong to $K$ and is not useful for our constructions. Second, the definition must be one-sided. Indeed, if one required the two-sided condition $\phi_{j}/\alpha\leqslant\theta_{j}\leqslant\alpha\phi_{j}$, then even in dimension $d=1$ the set $K=[1,L]$ with $L\geqslant 2$ would require at least $\log_{\alpha}L$ points. Third, convexity is required: for $d=2$, the nonconvex set
@@
also requires at least $\log_{\alpha}L$ points for any fixed $\alpha$. If desired, one could drop convexity but then relax the requirement to $A\subset\mathop{\mathrm{conv}}(K)$; we note, however, that this relaxation will not be useful for the applications discussed in this paper. Finally, even for convex $K$, an exponential dependence on $d$ is sometimes unavoidable for constant $\alpha$ (e.g., $K=\Delta_{d-1}$, the simplex in $\mathbb{R}^{d}$). A simple illustration is $\Delta_{1}=\{(t,1-t):t\in[0,1]\}\subset\mathbb{R}_{+}^{2}$: if $\alpha<2$ (say $\alpha=1.9$), no single point $\phi=(a,1-a)\in\Delta_{1}$ can $\alpha$-ratio cover both extreme points $(1,0)$ and $(0,1)$. Hence any $1.9$-ratio cover of $\Delta_{1}$ must have size at least $2$. Consequently, if $d$ is even, a $1.9$-ratio cover for the product set $\prod_{i=1}^{d/2}\Delta_{1}\subset\mathbb{R}_{+}^{d}$ must have size at least $2^{d/2}$: each block $\Delta_{1}$ contributes an independent binary choice, so one needs a distinct cover point for each of the $2^{d/2}$ combinations.
We provide two different approaches to the proof of Theorem˜1. The first is somewhat longer, but fully constructive and deterministic, based on first principles and exploiting the convexity of the set. The second proof, presented in Section˜4.3, is shorter and relies on a reduction to $\varepsilon$-net arguments presented in Diakonikolas and Yannakakis (2010) together with the volumetric argument used in the proof of Theorem˜3.
Proof roadmap for Theorem˜1.

### v2

@@ 1.5 Structure of the paper
In Section˜2 we study model aggregation over a finite family of densities and prove Theorem˜1. In Section˜3, we establish a sharp local covering bound in Hellinger distance for mixtures (Theorem˜2). Section˜4 is devoted to ratio covers, including the proof of Theorem˜3 and additional results on approximate Pareto curves. These tools are combined in Section˜5 to obtain the high-probability bound for convex aggregation (Theorem˜6). Further applications of ratio covers are presented in Section˜6: expected KL guarantees via a refined Yang–Barron argument (Section˜6.1) and KL covering numbers (Section˜6.2). We conclude with the misspecified setting in Section˜7.
@@ 2 Model aggregation
In this section, we provide a complete analysis of model aggregation over $M$ individual densities, where no assumptions are made on the individual densities. Despite its novelty, this result is technically less involved than the mixture case and can be seen as a preliminary building block introducing ideas that will also appear in this more complex setting. In particular, even at the preliminary stage, we do not need to use the Birgé-Le Cam tournaments, but rather exploit simple properties of the MLE in this case.
@@
This bound is optimal, in the sense that there are $M$ distributions such that no estimator can perform better up to universal constant factors, as shown in (Mourtada, 2025, Lemma 1) and (van der Hoeven et al., 2025, Theorem 11).
@@
  • Use the first $n$ observations (defining $\widehat{L}$) to construct a data-dependent set of almost empirical risk minimizers
    $\widehat{\mathcal{Q}}_{1}=\left\{p\in\mathcal{P}:\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\frac{2\log(4M/\delta)}{n}\right\}.$
@@
    $\widehat{\mathcal{Q}}_{2}=\bigg\{\frac{1}{2}p+\frac{1}{2|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q:p\in\widehat{\mathcal{Q}}_{1}\bigg\}.$ (8)
  • Using the second half of the sample, run the density estimator of Lemma˜2 over $\widehat{\mathcal{Q}}_{2}$ to output the final density $\widehat{p}$. Finally, if $75\log(2M)\log(2/\delta)\geqslant 2n\log(M)$ output instead $\widehat{p}=\frac{1}{M}\sum\limits_{p\in\mathcal{P}}p$.
Let us quickly explain why the procedure works. We first show that the construction of the set $\widehat{\mathcal{Q}}_{1}$ of almost empirical likelihood maximizers provides a set of small Hellinger radius that contains $p^{\star}$. However, closeness in Hellinger radius does not imply closeness in the Kullback–Leibler divergence. This is where the smoothed class $\widehat{\mathcal{Q}}_{2}$ helps: the construction does not inflate the Hellinger distance, but it allows one to control the density ratio. However, as an inevitable price of smoothing $p^{\star}$ does not have to belong to $\widehat{\mathcal{Q}}_{2}$. This is where now the regret-based techniques of Lemma 2 help, since they provide both the optimal tail bounds and allow to bypass the problem $p^{\star}\notin\widehat{\mathcal{Q}}_{2}$. Remarkably, the final estimator is optimal up to multiplicative constant factors.
@@
$\widehat{\mathcal{Q}}_{1}=\bigg\{p\in\mathcal{P}:\ \widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\frac{2\log(2M/\delta)}{n}\bigg\}.$
Then, with probability at least $1-\delta$, one has $p^{\star}\in\widehat{\mathcal{Q}}_{1}$ and
$\max_{p\in\widehat{\mathcal{Q}}_{1}}\operatorname{H}^{2}(p^{\star},p)\leqslant\frac{2\log(2M/\delta)}{n}.$
@@
$\mathbf{P}\left(\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant t\right)$ $=\mathbf{P}\left(\sqrt{\frac{\prod_{i=1}^{n}p(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}}\geqslant\exp(-nt/2)\right)$
@@
$\mathbf{P}\left(\exists p\in\mathcal{P}_{t}:\ \widehat{L}(p)-\widehat{L}(p^{\star})\leqslant t\right)\leqslant\sum_{p\in\mathcal{P}_{t}}\exp(-nt/2)\leqslant M\,\exp(-nt/2).$
Note that for any $p\in\mathcal{P}$, $\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant t$ implies $\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant t$, since $\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\widehat{L}(p^{\star})$. Therefore,
$\mathbf{P}\left(\exists p\in\mathcal{P}:\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant t\ \text{and}\ \operatorname{H}^{2}(p^{\star},p)\geqslant t\right)\leqslant M\exp(-nt/2).$
@@
$\mathbf{P}\left(\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant-t\right)$ $=\mathbf{P}\left(\sqrt{\frac{\prod_{i=1}^{n}p(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}}\geqslant\exp(nt/2)\right)$
@@
$\mathbf{P}\left(\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\widehat{L}(p^{\star})-t\right)\leqslant\sum_{p\in\mathcal{P}}\mathbf{P}\left(\widehat{L}(p)-\widehat{L}(p^{\star})\leqslant-t\right)\leqslant M\exp(-nt/2).$
@@
$\mathbf{P}\left(\widehat{L}(p^{\star})-\min_{q\in\mathcal{P}}\widehat{L}(q)\geqslant t\right)\leqslant M\exp(-nt/2).$
Choose $t={2\log(2M/\delta)}/{n}$. Then each of the two bad events above has probability at most $\delta/2$, so by a union bound they fail simultaneously with probability at most $\delta$. Hence, with probability at least $1-\delta$, every $p\in\widehat{\mathcal{Q}}_{1}$ satisfies $\operatorname{H}^{2}(p^{\star},p)\leqslant t$, and also $\widehat{L}(p^{\star})-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant t$ hence $p^{\star}\in\widehat{\mathcal{Q}}_{1}$. This yields the claim. ∎
@@
$\widehat{\mathcal{Q}}_{1}=\left\{p\in\mathcal{P}:\widehat{L}(p)-\min_{q\in\mathcal{P}}\widehat{L}(q)\leqslant\frac{2\log(4M/\delta)}{n}\right\}$
contains $p^{\star}$ and has the Hellinger diameter at most $2\sqrt{{2\log(4M/\delta)}/{n}}$. From now on we work on this event and focus on the second half of the sample. Using the second half of the sample we output the density $\widehat{p}$ obtained using the procedure of Lemma˜2 with respect to $\widehat{\mathcal{Q}}_{2}$. Let
$\widetilde{p}=\frac{1}{2}p^{\star}+\frac{1}{2|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q.$
@@
$\sup_{p,q\in\widehat{\mathcal{Q}}_{2}}\sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\leqslant\log\bigl(1+|\widehat{\mathcal{Q}}_{1}|\bigr)\leqslant\log(2M).$
@@
$\operatorname{KL}({p^{\star}},{\widehat{p}})\leqslant\operatorname{KL}({p^{\star}},{\widetilde{p}})+\frac{2\log(M)+\frac{25}{4}\left(e-2\right)\left(\log(2M)+1\right)\log\left({2}/{\delta}\right)}{n}.$ (9)
Finally, we have by Lemma 1 and the convexity of the Hellinger distance squared, observing that the ratio of $p^{\star}$ and $\frac{1}{2}p^{\star}+\frac{1}{2|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q$ is bounded by $2$,
@@
$\leqslant\frac{1}{(\sqrt{e}-1)^{2}}\operatorname{H}^{2}\left(p^{\star},\frac{1}{|\widehat{\mathcal{Q}}_{1}|}\sum\limits_{q\in\widehat{\mathcal{Q}}_{1}}q\right)\leqslant\frac{2\log(4M/\delta)}{(\sqrt{e}-1)^{2}n}.$
@@
Note that we used a sample of size $2n$; for any original sample size at least $2$, splitting into two equal parts (dropping at most one observation) increases the absolute constant by at most a factor of $3$ (the worst case is a total sample size of $3$). Finally, if $75\log(2M)\log(2/\delta)\geqslant 2n\log(M)$, we output $\widehat{p}=\frac{1}{M}\sum\limits_{p\in\mathcal{P}}p$, which guarantees $\operatorname{KL}(p^{\star},\widehat{p})\leqslant\log(M)$. The claim follows. ∎
@@ 4 Ratio covers of convex sets
After approximating the target distribution in Hellinger distance (as we did above, for instance in the special case of finite mixtures), a central question in this work is how to convert this control into a small set of candidate distributions such that at least one is close to $p^{\star}$ in KL divergence. We recall from Lemma 1 that the KL divergence $\operatorname{KL}({p},{q})$ can be bounded in terms of the Hellinger distance together with a factor depending on the maximal density ratio $\sup_{x\in\mathcal{X}}\frac{p(x)}{q(x)}$. In particular, if $p$ and $q$ lie in the same squared Hellinger ball of radius $\varepsilon$ and the density ratio is bounded by a constant, then $\operatorname{KL}({p},{q})\leqslant C\,\varepsilon$ for some constant $C>0$ depending only on that ratio bound. This motivates the general notion of a ratio cover, which we will later apply to Hellinger balls to obtain the KL covers needed in our applications.
@@
Our goal is to show the existence of $C_{1}$-ratio covers of size $(C_{2})^{d}$ for convex sets $K$, with constants $C_{1},C_{2}\geqslant 1$. In our proof of Theorem 6 for mixture models, we will ultimately pay a term proportional to $\log(|A|)/n$ for a ratio cover of size $|A|$: hence even a ratio cover of size $|A|=(\log d)^{d}$ would yield a suboptimal upper bound featuring an additional $\log\log d$ factor, and similarly any super-constant ratio $\alpha$ would yield suboptimal results (both issues are also important for our other applications). With this motivation, we state one of our main results.
Theorem 3.
@@
First, the constraint $A\subset K$ is essential: without it, there is a trivial “cover” obtained by choosing the single vector $\phi\in\mathbb{R}_{+}^{d}$ with coordinates $\phi_{j}=\max_{\theta\in K}\theta_{j}$, which need not belong to $K$ and is not useful for our applications. Second, the definition must be one-sided. Indeed, if one required the two-sided condition $\phi_{j}/\alpha\leqslant\theta_{j}\leqslant\alpha\phi_{j}$, then even in dimension $d=1$ the set $K=[1,L]$ with $L\geqslant 2$ would require at least $\log_{\alpha}L$ points. Third, convexity is required: for $d=2$, the nonconvex set
@@
also requires at least $\log_{\alpha}L$ points for any fixed $\alpha$. (Of course, one could drop the convexity assumption on $K$ but then relax the requirement to $A\subset\mathop{\mathrm{conv}}(K)$.) Finally, even for convex $K$, an exponential dependence on $d$ is sometimes unavoidable for constant $\alpha$. A simple illustration is $\Delta_{1}=\{(t,1-t):t\in[0,1]\}\subset\mathbb{R}_{+}^{2}$: if $\alpha=1.9$, no single point $\phi=(a,1-a)\in\Delta_{1}$ can $\alpha$-ratio cover both extreme points $(1,0)$ and $(0,1)$. Hence any $1.9$-ratio cover of $\Delta_{1}$ must have size at least $2$. Consequently, if $d$ is even, a $1.9$-ratio cover for the product set $\prod_{i=1}^{d/2}\Delta_{1}\subset\mathbb{R}_{+}^{d}$ must have size at least $2^{d/2}$: each block $\Delta_{1}$ contributes an independent binary choice, so one needs a distinct cover point for each of the $2^{d/2}$ combinations.
We provide two different approaches to the proof of Theorem˜3. The first is somewhat longer, but fully constructive and deterministic, based on first principles and exploiting the convexity of the set. The second proof, presented in Section˜4.3, is shorter and relies on a reduction to $\varepsilon$-net arguments presented in Diakonikolas and Yannakakis (2010) together with the volumetric argument used in the proof of Theorem˜2.
Proof roadmap for Theorem˜3.